write the equation of a line parallel to 4x+5y=20 and passing through (2,1)
Answer by mananth(16949)

4 x + 5 y = 20
Find the slope of this line
make y the subject ( rewrite in slope intercept form)
5y=-4x+20
Divide by 5
y =- ( 4/ 5)x+4
Compare this equation with y=mx+b
slope m = - 4/ 5
The slope of a line parallel to the above line will be the same
The slope of the required line will be- 4/ 5
m= - 4/ 5 ,point ( 2 , 1 )
Find b by plugging the values of m & the point in y=mx+b
1=-4/5 * 2+b
1 +8/ 5 = b
b= 13/ 5
m= - 4/5
Plug value of the slope and b in y = mx +b
The required equation is y = - 4/5 x + 13/5

m.ananth@hotmail.ca